Write the log equation for these coordinates: y= logbase(x)

(0.167, -1)

(1,0)

(6,1)

(36,2)


1
Expert's answer
2021-05-12T06:50:20-0400

"y=log_{base}(x)"

(a)

"-1=log_{base}(0.167)"

"base^{-1}=0.167"

"\frac{1}{base}=0.167"

"base=\frac{1}{0.167}"

"base\approx5.98"

therefore the logarithmic equation is:

"y=log_{5.98}(x)"


(b)

"0=log_{base}(1)"

"base^0=1...There\space will\space be\space an\space infinite\space solution\space for\space the\space base"

so, the base can be any number except 1 and negative numbers

in general, we can write,

"y=log(X)"


(c)

"1=log_{base}(6)\\base^1=6\\base=6"

Therefore the logarithmic equation is:

"y=log_6(x)"


(d)

"2=log_{base}(36)\\base^2=36\\base=6"

Therefore the logarithmic equation is:

"y=log_6(x)"